OpenStudy (ashchu117):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent? The area of the base of the tank The area of the curved sides of the tank The volume of 3 of the same cone-shaped tanks The volume of 6 of the same cone-shaped tanks

OpenStudy (ashchu117):

hold up

OpenStudy (ashchu117):

it didnt copy right

OpenStudy (ashchu117):

heres a screen shot

jhonyy9 (jhonyy9):

so r = h -4 ft but you know that h = x - 3 so this mean that r = x -3 -4 so r = x - 7 hope this will help you like a first step

OpenStudy (ashchu117):

im lost

OpenStudy (isaiah.feynman):



OpenStudy (isaiah.feynman):